# AI forecasting of higher-order wave modes of spinning binary black hole mergers

### v1

@@ 4 Results and Discussion
In Figure 8, we provide visualizations of the attention weights derived from the two decoder attention modules. These visualizations aid in forming an intuitive understanding of how the model attends to different parts of the input sequences for a given input waveform and choice of attention head. We present visualizations of attention weight matrices for all other attention heads in A.
@@
As detailed in the model section, the attention weights are computed using the dot product of queries and keys, $QK^{T}$, which measures the similarity between them. The softmax function then transforms these similarities into a probability distribution, resulting in a matrix of attention weights with shape $\mathbb{R}^{n_{dec}\times n_{enc}}$ in the case of the cross-attention module or with shape $\mathbb{R}^{n_{dec}\times n_{dec}}$ in the case of the self-attention module. These weights are used to weigh the contribution of the values $V$ for the next representation. For visualization purposes, we transpose these attention matrices, as shown in Figure 8.
@@
In both cross- and self-attention plots, we include the corresponding input sequences alongside the x- and y-axes to help visualize the model’s focus. For self-attention, the plotted decoder input sequence represents both queries (x-axis) and keys (y-axis). For cross-attention, the plotted decoder input sequence corresponds to the queries (x-axis), while the encoder input sequence represents the keys (y-axis). Although the attention modules operate on embeddings or intermediate representations rather than the raw sequences shown, the plotted sequences provide a useful reference for interpreting the attention patterns. Note that each decoder timestep $t_{dec}$ in the cross-attention plot (i.e., each column in the transposed weight matrix corresponding to some decoder timestep $t_{dec}$) contains contextual information about how each encoder timestep relates to the current decoder timestep, which is needed for the transformer to make the $t_{dec}+1$ prediction in the decoder output. Due to the softmax operation, all values in each column sum up to 1.
@@
The self-attention heatmap illustrates how each position in the decoder sequence attends to all other positions in the same sequence. For the self-attention mechanism in the decoder, each timestep can attend to all timesteps up to and including itself, but not to future timesteps, due to the causal attention mask. As a consequence of this property, there are no nonzero elements below the diagonal of the transposed weight matrix. As exemplified in Figure 8, the diagonal elements frequently tend to dominate in these self-attention weight plots, indicating that local context is especially important to the self-attention mechanism when making predictions.
The cross-attention heatmap reveals the model’s focus on different parts of the encoder output sequence while decoding each position of the output sequence. The attention is more diffused than in the self-attention case, as the model is integrating information from a broader range of encoder timesteps to inform each prediction in the decoder output.
These results, along with those presented in A, provide insight into how the attention heads prioritize different timesteps in the input signal during the forecasting process. The patterns shown in these visualizations change as we consider different input signals, highlighting the model’s sensitivity to variations in the input features. Interactive visualizations for different input signals are available at our GitHub repository [39].
5 Summary and conclusions
We have introduced a physics-inspired transformer model capable of learning and forecasting the nonlinear dynamics of higher-order wave modes that describe the physics of quasi-circular, spinning, non-precessing binary black hole mergers.
We developed algorithms to train AI models using over 14 million waveforms with distributed training, harnessing 16 NVIDIA A100 GPUs in the Delta supercomputer. This approach reduced time-to-solution to 15 hours. Furthermore, we used distributed inference to quantify the performance of our transformer model using 4 NVIDIA H100 GPUs in the DeltaAI supercomputer, finding that over a test set of 840,000 waveforms, our transformer model predicts waveforms whose mean and median overlaps with ground truth signals are 0.996 and 0.997, respectively. These results indicate that the transformer model learned the physics of these sources to a high level of accuracy. There are some regions, however, in which the AI model may be improved in future work, namely: low mass-ratios with anti-aligned spins, and high mass-ratio binaries that describe head-on mergers, i.e., with inclination angles $\theta=\pi/2$. Nonetheless, our results indicate that these samples represent a small set, since only 7.15% of test cases have overlaps below 0.99.
The work presented here, in terms of AI models, distributed training, and inference, paves the way for future studies that may consider the signal manifold of higher-order wave modes of spinning, precessing binary black hole mergers. This high dimensional signal manifold serves as a sandbox to stress test the capabilities of physics-inspired AI models and exascale computing to learn complex, non-linear dynamics of compact binary sources. This study should be pursued in future work.
@@ Acknowledgements
Appendix A Attention head visualizations
Here we present visualizations of self-attention and cross-attention weights for all attention heads, except for those presented in Figure 8. These results indicate that attention heads gather information across the input time span of input signals to produce accurate predictions.
[…]
ıin 0, 1, 2, 3, 4, 5, 7, 8, 9
Figure 9: Visualization of decoder self-attention weights for all 10 attention heads, excluding the one already presented in Figure 8, and choice of input waveform.
[…]
in 0, 1, 2, 3, 4, 5, 6, 7, 9
Figure 10: Visualization of cross-attention weights for all attention heads, excluding the one already presented in Figure 8, and choice of input waveform.

### v2

@@ 3.7 Attention Plots
In Figure 12, we provide visualizations of the attention weights derived from the two decoder attention modules. While primarily qualitative, these visualizations help form an intuitive understanding of how the attention mechanism operates across different input waveforms and attention heads.
@@
As detailed in the model section, the attention weights are computed using the dot product of queries and keys, $QK^{T}$, which measures the similarity between them. The softmax function then transforms these similarities into a probability distribution, resulting in a matrix of attention weights with shape $\mathbb{R}^{n_{\textrm{dec}}\times n_{\textrm{enc}}}$ in the case of the cross-attention module, or with shape $\mathbb{R}^{n_{\textrm{dec}}\times n_{\textrm{dec}}}$ in the case of the self-attention module. These weights are used to weigh the contribution of the values $V$ for the next representation.
@@
In both cross- and self-attention plots, we include the corresponding input sequences alongside the x- and y-axes to help visualize the model’s focus. For self-attention, the plotted decoder input sequence represents both queries (x-axis) and keys (y-axis). For cross-attention, the plotted decoder input sequence corresponds to the queries (x-axis), while the encoder input sequence represents the keys (y-axis). Although the attention modules operate on embeddings or intermediate representations rather than the raw sequences shown, the plotted sequences provide a useful reference for interpreting the attention patterns. Note that each decoder timestep $t_{\textrm{dec}}$ in the cross-attention plot (i.e., each column in the transposed weight matrix corresponding to some decoder timestep $t_{\textrm{dec}}$) contains contextual information about how each encoder timestep relates to the current decoder timestep, which is needed for the transformer to make the $t_{\textrm{dec}}+1$ prediction in the decoder output. Due to the softmax operation, all values in each column sum up to 1.
@@
The self-attention heatmap illustrates how each position in the decoder sequence attends to all other positions in the same sequence. For the self-attention mechanism in the decoder, each timestep can attend to all timesteps up to and including itself, but not to future timesteps, due to the causal attention mask. As a consequence of this property, there are no nonzero elements below the diagonal of the transposed weight matrix. As exemplified in Figure 12, the diagonal elements frequently tend to dominate in these self-attention weight plots, indicating that local context is especially important to the self-attention mechanism when making predictions. The cross-attention heatmap reveals the model’s focus on different parts of the encoder output sequence while decoding each position of the output sequence. The attention is more diffused than in the self-attention case, as the model is integrating information from a broader range of encoder timesteps to inform each prediction in the decoder output.
Interactive visualizations of the attention patterns for selected waveforms are available at  [29] to support further qualitative inspection. However, we caution against overinterpreting these attention plots as a reflection of the model’s reasoning. The patterns tend to correlate closely with the amplitude structure of the input waveforms, showing little abstraction or parameter-dependent structure across examples. These findings are consistent with prior work indicating that attention weights do not necessarily provide meaningful explanations for model behavior[52]. Accordingly, we present these visualizations as diagnostic and instructive tools rather than interpretability claims.
4 Discussion
This work demonstrates that transformer-based models can serve as fast, accurate surrogates for gravitational waveforms with higher-order modes, enabling waveform forecasting from quasi-circular, spinning, non-precessing binary black hole mergers across the late inspiral, merger, and ringdown phases. The model was trained on over 14 million waveforms using distributed training across 16 NVIDIA A100 GPUs on the Delta supercomputer, reaching convergence within 15 hours. Inference on 840,000 test waveforms was conducted using a single NVIDIA V100 GPU on the HAL system, with an average latency of 18 ms per waveform.
[…]
To assess generalization beyond the surrogate training domain, we benchmarked the model against numerical relativity waveforms from the SXS catalog. Despite including systems outside the training domain and additional harmonic content, the model achieved a median overlap of 0.969, with low-overlap outliers scattered only around $\theta=\pi/2$, indicating strong performance even on high-fidelity, out-of-distribution data. The high data-to-parameter ratio (40,000:1) and strong NR generalization suggest the model avoids overfitting.
Further analysis revealed that the model relies heavily on late-inspiral information in systems with high spin and mass asymmetry, while earlier inspiral features proved more important for near-equal-mass binaries. The model also appears to extract inclination angle-dependent structure well before merger. Mode truncation experiments confirmed that higher-order modes significantly enhance accuracy, particularly in systems with large mass ratios and edge-on orientations.
[…]
This work provides a starting point for future research into higher-order wave modes, including those from spinning, precessing binary black hole mergers. This broader signal manifold offers an opportunity to systematically explore the capabilities of AI models and exascale computing in learning complex, nonlinear dynamics of compact binary systems. These directions should be pursued in future work to further deepen our understanding of gravitational wave signals.
@@ Acknowledgements
Appendix A Fine-Grained Analysis via Spin Slices
Figure 13: Reconstruction accuracy across mass ratio, $q$, and inclination angle, $\theta$, for several fixed values of the effective spin parameter, $s_{\mathrm{eff}}^{z}$.
Figure 13 shows the test set partitioned into fixed-spin slices across the $(q,\theta)$ plane for various values of effective spin, $s_{\mathrm{eff}}^{z}$. This stratified analysis reveals that the transformer model consistently achieves high predictive accuracy for binaries with strongly aligned or anti-aligned spins, i.e., $s_{\mathrm{eff}}^{z}\sim\pm 0.8$. Within these regimes, the model performs robustly across a wide range of mass ratios and inclination angles.
However, at intermediate spin values—particularly around $s_{\mathrm{eff}}^{z}\sim 0$—a localized degradation in performance is observed. Specifically, reduced overlap values are concentrated around systems with comparable mass ratios ($q\sim 1$) and face-on orientations ($\theta\sim 0$), suggesting a sensitivity of the model to configurations with low intrinsic spin alignment and minimal inclination-induced modulation.
As $s_{\mathrm{eff}}^{z}$ departs from zero in either direction, predictive accuracy systematically improves across the $(q,\theta)$ slices, highlighting the model’s effectiveness in capturing spin-modulated waveform features. Nonetheless, the model exhibits a consistent dip in accuracy for edge-on binaries ($\theta\sim\pi/2$), irrespective of spin, indicating a challenging regime associated with vanishing cross polarization and potential waveform degeneracies.
Finally, while the model attains near-optimal performance for $s_{\mathrm{eff}}^{z}\sim\pm 0.8$, we identify a narrow region of reduced accuracy at $s_{\mathrm{eff}}^{z}\sim-0.8$ for high-mass-ratio, edge-on systems ($q\sim 8$, $\theta\sim\pi/2$). This subtle degradation suggests a complex interplay between spin orientation, geometric projection effects, and mass asymmetry in these extreme regions of the parameter space.
Appendix B Binned Overlap Drop by Parameter for Obfuscation Study
To complement the correlation analysis in section 3.6.1, we provide binned heatmaps of the average overlap drop as a function of parameter value and obfuscated input range. These plots offer a finer-grained view of parameter influence on predictions, allowing us to assess possible nonlinear or symmetric effects that may be obscured by Pearson correlation.
Figure 14 shows match drop by mass ratio $q$. Most input segments exhibit a broadly monotonic increase in match drop with $q$, except in the $(-1000\textrm{M},-200\textrm{M})$ segment, where it decreases. Early segments show slight non-monotonicity, particularly near $q\approx 1$, explaining the low Pearson correlation in that region. Figure 15 displays match drop by primary spin $s_{1}^{z}$. The trend is mostly monotonic, with drop increasing with spin. A deviation appears for $s_{1}^{z}<-0.5$, where drop temporarily decreases. In the $(-1000\textrm{M},-200\textrm{M})$ segment, match drop instead decreases with spin.
Figure 14: Binned overlap drop by mass ratio $q$. Most input segments show a broadly monotonic relationship, with match drop increasing for more asymmetric binaries (with the exception of the $(-1000\textrm{M},-200\textrm{M})$ range, where match drop decreases monotonically with $q$). Some early segments display small deviations from this monotonicity, particularly at low $q$. Notably, the low Pearson correlation observed in the earliest segment is largely driven by this localized effect around $q\approx 1$.
Figure 15: Binned overlap drop by primary spin $s_{1}^{z}$. The overall trend is approximately monotonic across the spin range, with match drop generally increasing with spin (except in the $(-1000\textrm{M},-200\textrm{M})$ range, where match drop decreases with spin). A slight deviation from monotonicity is observed for very negative spin values $(s_{1}^{z}<-0.5)$, where match drop temporarily decreases before continuing the broader trend.
Figure 16 shows match drop by secondary spin $s_{2}^{z}$, which is relatively flat across all bins. Figure 17 plots match drop by inclination angle $\theta$. Most segments show a symmetric dependence around $\theta\approx\pi/2$, justifying the use of $|\theta-\pi/2|$ in the main analysis. In the final segment, this symmetry breaks slightly, with drop increasing toward higher inclination.
Figure 16: Binned overlap drop by secondary spin $s_{2}^{z}$. The overlap drop is relatively flat across all bins (i.e., within each column).
Figure 17: Binned overlap drop by inclination angle $\theta$. Most time segments show a roughly symmetric dependence around $\theta\approx\pi/2$, motivating the use of the transformed variable $\lvert\theta-\pi/2\rvert$ in the main analysis. In the final segment $(-200\textrm{M},-120\textrm{M})$, this symmetry breaks down slightly, with match drop increasing toward higher inclination angles.